# Generating Jackiw-Teitelboim Euclidean gravity from static three-dimensional Maxwell-Chern-Simons electromagnetism

### v1

@@ 5 Gauss-Bonnet theorem: computing sizes
The GB theorem [^6] In the discussion provided here of the GB theorem, we attain ourselves to two-dimensional Riemannian manifolds. For generalizations to any dimension, see [68, 69]. [63, 65] is a powerful tool in the study of manifolds. The standard GB theorem for a connected and compact manifold, say $M^{2}$, states that [63, 65],
@@
with $\chi$ being the Euler characteristic of the manifold. It is known that the GB theorem (5.1) generalizes for complete, finitely connected, and non-compact manifolds [70, 71, 72, 73] if the integral is absolutely integrable and the total area is finite. Nevertheless, we mostly confine ourselves to the compact surface cases.
Formula (5.1) can be employed to select more accurately the possible manifolds of our solutions. And, in some cases, it can be employed to determine their sizes, according to their Euler characteristics.
@@ 5.1 Flat surfaces
Solution encountered in Section 3 is a locally flat manifold of minimal area. Considering the GB theorem for compact surfaces of vanishing curvature, GB theorem in the form (5.1) restricts the pool of solutions to manifolds with zero Euler characteristic. It turns out that the manifold satisfying all requirements of the GB theorem is the torus with only one handle [63].
It turns out that formula (5.1) can also be applied for non-compact surfaces such as the cylinder [63, 74]. In fact, any manifold topologically equivalent to a cylinder or a torus, while maintaining the flatness of the surface, can be considered as a solution. For example, a disk with a hole in it satisfies the criteria. Observe that the cylinder and the holed disk are topologically equivalent; both are obtained from the continuous deformation of a sphere with two holes.
@@
Figure 1: Flat surfaces with vanishing Euler characteristics. From left to right: The cylinder; the flat disk with a hole; the flat torus. Among these manifolds, only the flat torus is compact.
@@ 5.2 Hyperbolic surfaces
The solution found in Section 4.1 is a manifold of negative constant curvature, i.e., any hyperbolic surface [63, 64]. Typical non-compact examples are the tractroid (pseudo-sphere) and the catenoid (See Figure (2)), which is homeomorphic to the cylinder, and thus has a vanishing Euler characteristic as well [63]. Therefore, we cannot employ the GB theorem (5.1) to compute its size. For compact hyperbolic surfaces, the simplest manifolds with negative curvature are essentially those which are equivalent to the $n$ linearly connected torii ($n$-holed torus), where $n$ is the number of handles (related to its genus) and the corresponding Euler characteristic is given by $\chi=2-2n$, with $n>1$ (The case $n=1$ is the simple torus with only one handle and vanishing curvature - see the previous Section) [63, 64]. Clearly, for hyperbolic surfaces, $\chi$ must be negative. Of course, the torii can be non-linearly glued, but the resulting genus would be bigger than $n$, as we will approach in a following example. Thus, putting the solution (4.8) in the GB theorem (5.1), we find [^7] We remark that it is commonly known the validity of the formula $\mathcal{A}_{h}=-2\pi\chi\;,$ (5.2) for the area of compact surfaces [63]. At a first look, this formula seems to not be compatible with formula (5.4). Nevertheless, this formula is deduced for surfaces with normalized curvature, which is not the case here due to the presence of the cosmological constant.
@@
Figure 2: Non-compact hyperbolic manifolds. From left to right: The tractroid; and the catenoid. These manifolds have vanishing Euler characteristics.
Let us consider a first example, for instance (See Figure 3), $n$ identical torii, linearly connected, with two equal radii $2b_{h}$ and $b_{h}$, and making the interface between each torii be a circle of radius $c<<b_{h}$ (In this way, we characterized the manifold by only one parameter). This is the well known $n$-holed torus. We can determine the radius $b_{h}$ from formula (5.5):
@@
The minimum possibility is thus $n=2$, providing $b_{h}=0.87\mu m$. As $n\longrightarrow\infty$, the area stabilizes at the maximum value of $b_{h}=1.23\mu m$. We conclude that, although the area diverges with $n$, the radii of the identical $n$-holed torus do not.
Figure 3: The $n$-holed torus. A compact hyperbolical surface with non-vanishing Euler characteristics ($\chi=2-2n$). To compute the size of this manifold in an easy way, the intersections of the torii are made very small.
Just to illustrate another case, where the genus changes, we consider now the same $n$-holed torus of the previous example. However, we glue the first and last torii together, in a kind of ring made by $n$ torii, see Figure 4. In this case, the genus increases by one, $\chi=2-2(n+1)=-2n$. Thence,
@@
Figure 4: The closed $n$-holed torus. They are linearly connected and the torii of the edges are also connected, forming a ring of torii. The big hole in the middle increases its Euler characteristics by one ($\chi=-2n$). Just like the open linearly connected $n$-holed torus, the intersections of the torii are made very small in order to compute the size of this manifold in an easy way.
These results are quite remarkable because it essentially says that the sizes of some surfaces in the present gravity model are determined ultimately by the electric charge in the original model, i.e., the manifold scale in one theory is related to the electric charge in another theory.
@@ 5.3 Spherical surfaces
In Section 4.2, another sort of solution was derived. Essentially, manifolds with constant positive curvature (see (4.14)). Traditional non-compact examples are spherical caps and curved disks. For compact solutions, we have the usual sphere itself.
@@
Among all possible surfaces, we take the standard sphere as an example, see Figure 5. Therefore, direct integration of formula (4.14) and the usage of expression (5.1) imply the relation $\Lambda^{2}\mathcal{A}_{s}=4\pi\chi$, with $\mathcal{A}_{s}$ being the area of the sphere. For the sphere, the Euler characteristic is given by $\chi=2$. As a consequence, its area is fixed by $\mathcal{A}_{s}=8\pi/\Lambda^{2}$. And the radius $b_{s}$ of the sphere is thus determined by the simple formula
@@
Substituting the numerical value of the cosmological constant, as computed in expression (3.5), we get
@@
Therefore, by considering as an example the system of $n$ identical torii of the previous Section, we get
@@
in complete accordance with the first example in the previous Section.
@@ 6 Conclusions and perspectives
In this paper, we have shown that pure three-dimensional electrodynamics at the static limit corresponds to two-dimensional dilatonic gravity in Euclidean space. The Newtonian and cosmological constants are completely determined by the electric charge, as inferred in the Relations (3.3). Moreover, we were able to compute these constants, see Equation (3.5). Both of them are found to be several orders of magnitude larger than their four-dimensional experimental values, implying a kind of strong two-dimensional gravity.
@@
We also demonstrated three possible solutions for such a gravity model in Sections 3 and 4, namely flat, hyperbolic, and spherical manifolds. These solutions essentially constitute the three possible situations of two-dimensional manifolds [63]. In Section 5, we employed the GB theorem to determine their sizes and some topological properties. Moreover, in all three solutions, the dilaton is a constant arbitrary field. A fact that is consistent with its electromagnetic origin as the scalar potential. The main properties of the solutions discussed in these Sections are described in what follows:
@@
  • Flat manifolds: This solution is obtained from the action (3.4), originating directly from the mapping. In two dimensions, for several types of manifolds, the GB theorem, in formula (5.1), can be used to integrate the curvature scalar. Among these manifolds, since the curvature vanishes, we are restricted to manifolds with Euler characteristics zero. The possibilities we have inferred as examples are the cylinder (or, equivalently, a flat disk with a hole in it) and the flat torus (See Figure 1 for the examples). The flat torus is, of course, the only compact surface in these examples. Unfortunately, the size of these manifolds cannot be determined by the GB theorem.
@@
  • Hyperbolic manifold: This solution is obtained by enforcing the shift constraint (4.3), related to the shift (4.2). The result is a JT-type of gravity with manifolds of constant negative curvature. Among hyperbolic surfaces, the tractroid, the catenoid, and the $n$-holed torus (See Figure 2 for the non-compact surfaces and Figures 3 and 4 for the compact examples) can be taken as examples. Since the catenoid and the tractroid are topologically equivalent to the cylinder, they have vanishing Euler characteristics, disabling the usage of the GB theorem (5.1) to determine the sizes of these surfaces.
@@
    For the $n$-holed torii, on the other hand, we have applied the GB theorem in order to find their sizes in two examples: the open linearly connected $n$-holed torus (See Figure 3) and the closed linearly connected $n$-holed torus (Figure 4). The area of these manifolds depend on $n$ and $\Lambda$, as demonstrated in Expressions (5.3) and (5.7). Explicit computations were performed by considering a specific size of each torus, taken to be identical and characterized only by one parameter, the radius of the inner circle of the transversal cross-section of one torus. It was found an incredibly small radius for these manifolds. For the open $n$-holed torus with $n=2$, for example, the radius is just $0.87\mu m$. As $n$ increases, the radius asymptotically stabilizes at the value of $1.23\mu m$. For the closed $n$-holed torus, the radius does not depend on $n$ and is fixed to the value of $1.23\mu m$. Obviously, these values are computed from the value of the cosmological constant, calculated in (3.5), originally determined by the electric charge.
@@
  • Spherical manifolds: This solution is obtained via the shift (4.11) and the (same as the hyperbolic solution) shift constraint (4.3). Constant positive curvature for JT gravity is obtained, namely (4.14). Therefore, the solution describes spherical manifolds. Among spherical surfaces, we pick as example the usual sphere (Figure 5), which is compact. Thence, GB theorem can easily be applied in order to determine the sphere radius in terms of the cosmological constant. Following the same steps developed in the $n$-holed torii examples, we were able to find that the radius of the sphere is astonishingly small, valuing $3.09\mu m$.
@@
A more technical comment is that the constraint (4.3) (or the corresponding shifts of the curvature and spin-connection) is, in practice, equivalent to a choice of a background configuration. Particularly, hyperbolical or spherical ones. This background can only be chosen due to the presence of mass parameters in the action. Otherwise, a mass parameter would be required by hand in order to achieve JT gravity. Fortunately, we just do not have any mass parameter, but the specific parameter interpreted as a cosmological constant. Even if we did not have it, we recall that $G$ also carries mass dimension, allowing a different choice of curvature values.
@@
A remarkable feature of the model is that the relation between the gravitational constants and the electric charge, given in (3.3), indicates a possible duality because $G\sim 1/e^{2}$. Thence, if on one hand we have electrodynamics as a perturbative system, on the other hand we have a JT gravity model with a strong coupling parameter. Alternatively, there is no obstruction in starting from gravity with any value of $G$ and map it in static MCS theory while maintaining the relation between $G$ and $e$. In that case, one could describe standard JT gravity in terms of MCS static electromagnetism with some exotic value of electric charge. Thus, for example, a black hole configuration in JT gravity could be described by an electromagnetic system with a specific kind of charged fields. Such ideas also suggest that our model also has the potential to be employed in condensed matter systems, as discussed in the Introduction.
@@
A few extra comments are in order. First, it is important to emphasize that we have used a map between pure static three-dimensional MCS electromagnetism and two-dimensional Euclidean gravity in the FOF. The electromagnetism side is a generic model with no characterized length scale (or energy scale). The only number we have considered is the fundamental charge of the electron $e$. On the gravity side, we found three possible solutions. Two of them have specific length scales of the order of $\mu m$ (micrometers). This feature originates from the fact that the curvatures of the exemplified surfaces are determined from the cosmological constant, while the latter is fixed by the electric charge. In essence, these two solutions are solutions of JT gravity with negative and positive curvatures.
@@
In the same spirit, we can start in the gravity model we found and go back to electrodynamics at the static limit. Therefore, one could think of associating systems with characteristic lengths of the order of $\mu m$ with the properties of static MSC electrodynamics. Examples in Nature of objects and living beings of such small size are: suspended fine dust; the mitochondria; large viruses such as the Megavirus chilensis [75] and the Acanthamoeba polyphaga Mimivirus [76, 77]; some bacteria (e.g., the Escherichia coli, also known as the E. coli, [78]). Thus, perhaps, the surface of these systems could be mimicked by MCS static theory in flat space.
@@
In truth, there are many perspectives for future investigations. We mention here some immediate possibilities. For instance, we can look at the four-dimensional case and its corresponding dimensional reduction (time and one space coordinate, for example). Naturally, one extra scalar field will emerge, and the dynamics can change. Another point is the inclusion of matter. Fermions could be considered in the same way we have considered them in [56], or one could simply include external source interactions. Of course, extra care will be demanded in the dimensional reduction of spinors [79, 80, 81]. It could also be interesting to study matter with different electric charges. In this case, the violation of the equivalence principle is expected since only one type of charge can be absorbed in the gauge field. Moreover, inclusion of matter may be relevant for black hole analysis [24, 26, 5]. In terms of time evolution, one can simply decompose space and time of three-dimensional electrodynamics and map to gravity only the space coordinates. The expected result would be that of surfaces and the dilaton evolving in time as a separate coordinate of the manifold. The same idea can be generalized to four dimensions with only one space coordinate dimensional reduction - in that case, two scalar fields will appear, possibly evolving in time together with the manifold.
@@
Finally, it is worth mentioning that, in two dimensions, the map (3.1) is also consistent at the quantum level. Therefore, one can expect that the same feature is valid in the present case. Nevertheless, this study is beyond the scope of the present work.

### v2

@@ 5 Gauss-Bonnet theorem: computing sizes
The GB theorem [^6] In the discussion provided here of the GB theorem, we restrict ourselves to two-dimensional Riemannian manifolds. For generalizations to any dimension, see [69, 70]. [64, 66] is a powerful tool in the study of manifolds. The standard GB theorem for a connected and compact manifold, say $M^{2}$, states that [64, 66],
@@
with $\chi$ being the Euler characteristic of the manifold. It is known that the GB theorem (5.1) generalizes for complete, finitely connected, and noncompact manifolds [71, 72, 73, 74] if the integral is absolutely integrable and the total area is finite. Nevertheless, we mostly confine ourselves to the compact surface cases.
The formula (5.1) can be employed to select more accurately the possible manifolds of our solutions. And, in some cases, it can be employed to determine their sizes, based on their Euler characteristics.
@@ 5.1 Flat surfaces
The solution discussed in Section 3 is a locally flat manifold of minimal area. Considering the GB theorem for compact surfaces of vanishing curvature, the GB theorem in the form (5.1) restricts the pool of solutions to manifolds with zero Euler characteristic. It turns out that the manifold satisfying all requirements of the GB theorem is the torus with only one handle [64].
It turns out that formula (5.1) can also be applied to noncompact surfaces such as the cylinder [64, 75]. In fact, any manifold topologically equivalent to a cylinder or a torus, while maintaining the flatness of the surface, can be considered as a solution. For example, a disk with a hole in it satisfies the criteria. Notice that the cylinder and the holed disk are topologically equivalent; both are obtained from the continuous deformation of a sphere with two holes.
@@
Figure 1: Flat surfaces with vanishing Euler characteristic. From left to right: The cylinder; the flat disk with a hole; the flat torus. Among these manifolds, only the flat torus is compact.
@@ 5.2 Hyperbolic surfaces
The solution found in Section 4.1 is a manifold of negative constant curvature, i.e., any hyperbolic surface [64, 65]. Typical noncompact examples are the tractroid (pseudo-sphere) and the catenoid (See Figure (2)), which is homeomorphic to the cylinder, and thus has a vanishing Euler characteristic as well [64]. Therefore, we cannot employ the GB theorem (5.1) to compute its size. For compact hyperbolic surfaces, the simplest manifolds with negative curvature are essentially those which are equivalent to the $n$ linearly connected tori ($n$-holed torus), where $n$ is the number of handles (related to its genus) and the corresponding Euler characteristic is given by $\chi=2-2n$, with $n>1$ (The case $n=1$ is the simple torus with only one handle and vanishing curvature - see the previous Section) [64, 65]. Clearly, for hyperbolic surfaces, $\chi$ must be negative. Of course, the tori can be nonlinearly glued, but the resulting genus would be bigger than $n$, as we will approach in a following example. Thus, putting the solution (4.8) in the GB theorem (5.1), we find [^7] We remark that it is commonly known the validity of the formula $\mathcal{A}_{h}=-2\pi\chi\;,$ (5.2) for the area of compact surfaces [64]. At a first glance, this formula seems to not be compatible with formula (5.4). Nevertheless, this formula is deduced for surfaces with normalized curvature, which is not the case here because of the presence of the cosmological constant.
@@
Figure 2: Noncompact hyperbolic manifolds. From left to right: The tractroid; and the catenoid. These manifolds have vanishing Euler characteristics.
Let us consider a first example, for instance (see Figure 3), $n$ identical tori, linearly connected, with two equal radii $2b_{h}$ and $b_{h}$, and making the interface between each tori a circle of radius $c<<b_{h}$ (In this way, we characterized the manifold by only one parameter). This is the well-known $n$-holed torus. We can determine the radius $b_{h}$ from formula (5.5):
@@
The minimum possible value is thus $n=2$, providing $b_{h}=0.87\mu m$. As $n\longrightarrow\infty$, the area stabilizes at the maximum value of $b_{h}=1.23\mu m$. We conclude that, although the area diverges with $n$, the radii of the identical $n$-holed torus do not.
Figure 3: The $n$-holed torus. A compact hyperbolical surface with non-vanishing Euler characteristics ($\chi=2-2n$). To compute the size of this manifold in an easy way, the intersections of the tori are made very small.
Just to illustrate another case, where the genus changes, we consider now the same $n$-holed torus of the previous example. However, we glue the first and last tori together, to form ring composed of $n$ tori, see Figure 4. In this case, the genus increases by one, $\chi=2-2(n+1)=-2n$. Thence,
@@
Figure 4: The closed $n$-holed torus. They are linearly connected and the tori of the edges are also connected, forming a ring of tori. The big hole in the middle increases its Euler characteristics by one ($\chi=-2n$). Just like the open linearly connected $n$-holed torus, the intersections of the tori are made very small in order to compute the size of this manifold in an easy way.
These results are quite remarkable because it essentially says that the sizes of some surfaces in the present gravity model are ultimately determined by the electric charge in the original model, i.e., the manifold scale in one theory is linked to the electric charge in another theory.
@@ 5.3 Spherical surfaces
In Section 4.2, another sort of solution was derived. Essentially, manifolds with constant positive curvature (see (4.14)). Traditional noncompact examples include spherical caps and curved disks. For compact solutions, we have the usual sphere itself.
@@
Among all possible surfaces, we consider the standard sphere as an example (see Figure 5). Therefore, direct integration of formula (4.14) and the use of expression (5.1) imply the relation $\Lambda^{2}\mathcal{A}_{s}=4\pi\chi$, with $\mathcal{A}_{s}$ being the area of the sphere. For the sphere, the Euler characteristic is given by $\chi=2$. As a consequence, its area is fixed by $\mathcal{A}_{s}=8\pi/\Lambda^{2}$ and the radius $b_{s}$ of the sphere is thus determined by the simple formula
@@
Substituting the numerical value of the cosmological constant, as computed in expression (3.5), we obtain
@@
Therefore, considering as an example the system of $n$ identical tori from the previous Section, we get
@@
in full agreement with the first example in the previous Section.
@@ 6 Conclusions and perspectives
In this paper, we have shown that pure three-dimensional electrodynamics in the static limit corresponds to two-dimensional dilatonic gravity in Euclidean space. The Newtonian and cosmological constants are completely determined by the electric charge, as indicated in the relations (3.3). Furthermore, we were able to compute these constants, see Equation (3.5). Both of them are found to be several orders of magnitude larger than their four-dimensional experimental values, indicating a form of strong two-dimensional gravity.
@@
We also presented three possible solutions for such a gravity model in Sections 3 and 4, namely flat, hyperbolic, and spherical manifolds. These solutions essentially constitute the three possible situations of two-dimensional manifolds [64]. In Section 5, we employed the GB theorem to determine their sizes and some topological properties. Moreover, in all three solutions, the dilaton is a constant arbitrary field, which is consistent with its electromagnetic origin as a scalar potential. The main properties of the solutions discussed in these Sections are described in what follows:
@@
  • Flat manifolds: This solution is obtained from the action (3.4), which originates directly from the mapping. In two dimensions, for various types of manifolds, the GB theorem, in formula (5.1), can be used to integrate the scalar curvature. Among these manifolds, since the curvature vanishes, we are restricted to manifolds with zero Euler characteristic. The possibilities we have inferred as examples are the cylinder (or, equivalently, a flat disk with a hole in it) and the flat torus (see Figure 1 for the examples). The flat torus is, of course, the only compact surface in these examples. Unfortunately, the size of these manifolds cannot be determined by the GB theorem.
@@
  • Hyperbolic manifold: This solution is obtained by enforcing the shift constraint (4.3), related to the shift (4.2). The result is a JT-type gravity with manifolds of constant negative curvature. Among hyperbolic surfaces, the tractroid, the catenoid, and the $n$-holed torus (see Figure 2 for the noncompact surfaces and Figures 3 and 4 for the compact examples) can be taken as examples. Since the catenoid and the tractroid are topologically equivalent to the cylinder, they have vanishing Euler characteristics, preventing the use of the GB theorem (5.1) to determine the sizes of these surfaces.
@@
    For the $n$-holed tori, we have applied the GB theorem to determine their sizes in two examples: the open linearly connected $n$-holed torus (see Figure 3) and the closed linearly connected $n$-holed torus (Figure 4). The area of these manifolds depend on $n$ and $\Lambda$, as shown in expressions (5.3) and (5.7). Explicit computations were performed by considering a specific size of each torus, taken to be identical and characterized only by one parameter, the radius of the inner circle of the transversal cross-section of one torus. It was found an incredibly small radius for these manifolds. For the open $n$-holed torus with $n=2$, for example, the radius is just $0.87\mu m$. As $n$ increases, the radius asymptotically stabilizes at the value of $1.23\mu m$. For the closed $n$-holed torus, the radius does not depend on $n$ and is fixed to the value of $1.23\mu m$. Naturally, these values are computed from the value of the cosmological constant, determined in (3.5), originally determined by the electric charge.
@@
  • Spherical manifolds: This solution is obtained through the shift (4.11) and the shift constraint (4.3), which is the same as for the hyperbolic solution. Constant positive curvature for JT gravity is obtained, namely (4.14). Therefore, the solution describes spherical manifolds. Among spherical surfaces, we take the usual sphere (Figure 5), which is compact, as an example. Thence, GB theorem can be readily applied in order to determine the sphere radius in terms of the cosmological constant. Following the same steps developed in the $n$-holed tori examples, we were able to find that the radius of the sphere is remarkably small, valuing $3.09\mu m$.
@@
A more technical comment is that the constraint (4.3) (or the corresponding shifts of the curvature and spin connection) is, in practice, equivalent to choosing a background configuration. Particularly, hyperbolical or spherical ones. This background can only be chosen due to the presence of mass parameters in the action. Otherwise, a mass parameter would be required by hand in order to achieve JT gravity. Fortunately, we do not have any such mass parameter, but the specific parameter interpreted as a cosmological constant. Even if we did not have it, we recall that $G$ also carries mass dimension, allowing a different choice of curvature values.
@@
A remarkable feature of the model is that the relationship between the gravitational constants and the electric charge, given in (3.3), indicates a possible duality because $G\sim 1/e^{2}$. Thence, if on one hand we have electrodynamics as a perturbative system, on the other hand we have a JT gravity model with a strong coupling parameter. Alternatively, there is no obstruction in starting from gravity with any value of $G$ and map it in static MCS theory while maintaining the relation between $G$ and $e$. In that case, one could describe standard JT gravity in terms of MCS static electromagnetism with some exotic value of electric charge. Thus, for example, a black hole configuration in JT gravity could be described by an electromagnetic system with a specific kind of charged fields. Such ideas also suggest that our model also has the potential to be employed in condensed matter systems, as discussed in the Introduction.
@@
A few additional remarks are in order. First, it is important to emphasize that we have used a map between pure static three-dimensional MCS electromagnetism and two-dimensional Euclidean gravity in the FOF. The electromagnetism side is a generic model with no characterized length scale (or energy scale). The only number we have considered is the fundamental charge of the electron $e$. On the gravity side, we found three possible solutions. Two of them have specific length scales of the order of $\mu m$ (micrometers). This feature originates from the fact that the curvatures of the exemplified surfaces are determined from the cosmological constant, while the latter is fixed by the electric charge. In essence, these two solutions are solutions of JT gravity with negative and positive curvatures.
@@
In the same spirit, we can start from the gravity model we found and go back to electrodynamics at the static limit. Therefore, one might consider associating systems with characteristic lengths on the order of $\mu m$ with the properties of static MSC electrodynamics. Examples in nature of objects and living beings of such small size are: suspended fine dust; the mitochondrion; large viruses such as the Megavirus chilensis [76] and the Acanthamoeba polyphaga Mimivirus [77, 78]; some bacteria (e.g., the Escherichia coli, also known as the E. coli, [79]). Thus, perhaps, the surface of these systems could be modeled by MCS static theory in flat space.
@@
Indeed, there are many perspectives for future investigations. We mention here some immediate possibilities. For instance, we can look at the four-dimensional case and its corresponding dimensional reduction (time and one space coordinate, for example). Naturally, one extra scalar field will emerge, and the dynamics can change. Another point is the inclusion of matter. Fermions could be considered in the same way we have considered them in [57], or one could simply include external source interactions. Of course, extra care will be demanded in the dimensional reduction of spinors [80, 81, 82]. It could also be interesting to study matter with different electric charges. In this case, the violation of the equivalence principle is expected since only one type of charge can be absorbed in the gauge field. Moreover, inclusion of matter may be relevant for black hole analysis [26, 28, 6]. In terms of time evolution, one can simply decompose space and time of three-dimensional electrodynamics and map to gravity only the space coordinates. The expected result would be that of surfaces and the dilaton evolving in time as a separate coordinate of the manifold. The same idea can be generalized to four dimensions with only one space coordinate dimensional reduction - in that case, two scalar fields will appear, possibly evolving in time together with the manifold. One last perspective to address is that one could extend the discussion developed here to the case of general affine connections with non-metricity. Such work would require a separate and more detailed investigation, which goes beyond the scope of the present paper.
@@
Finally, it is worth mentioning that, in two dimensions, the map (3.1) is also consistent at the quantum level. Therefore, one can expect that the same feature to hold in the present case. Nevertheless, this study is beyond the scope of the present work.